An electric dipole with dipole moment of magnitude 3.75 x 28 10 Cm is in an electric field of magnitude 9500 N/C. Compute the magnitude of the torque on the dipole when the dipole moment is (a) perpendicular to the electric field, (b) parallel to the electric field and (c) antiparallel to the electric field.


1
Expert's answer
2022-01-13T09:22:39-0500

Answer

Given dat

Electric field

"P=3.75\times 28.10Cm"

electric field of magnitude

E=9500 N/C

Now torque is gien by

"\tau =PE \sin\theta"

a) for parallel "\theta =0\\ \sin0°=0\\"

Torque

"\tau=0"

b) for antiparallel

"\theta=180°\\ \sin180°=0"

So torue

"\tau=0"